x^3-x^2-12x=0
This is a summer assignment that will be a major test grade when we go back to school in August. Thank you!!

Answer by checkley75(3666)

X^3-X^2-12X=0
X(X^2-X-12)=0
X(X-4)(X+3)=0
X=0 ANSWER
X-4=0 OR X=4 ANSWER.
X+3=0 OR X=-3 ANSWER.